Our examples in the chapter have focused on manufacturing, where the output is units of product and the inputs are manufacturing activities or costs. The concept of productivity can be applied in a variety of settings, wherever there are inputs and outputs. For example, consider LandscapeCity, a landscape design company that specializes in small landscape projects for people living in cities. Amanda Caldwell, the assistant manager, is in the process of trying to determine if productivity has been improving since she was hired 6 months ago. Because it is a design firm, labor is the only significant expense, but Amanda is unsure if the number of projects or the dollar of sales volume should be used when computing productivity. She has collected these data for sales and labor expenses for the past 6 months:

Month Labor Expense Number of Projects Sales Dollars
1 $ 21,310 44 $ 21,090
2 28,800 60 22,900
3 20,392 60 29,780
4 14,194 15 19,560
5 14,926 20 22,680
6 21,592 60 22,520

Required:

1. Calculate the productivity for each month and the change in productivity from month to month using number of projects as the measure of output.

2. Calculate the productivity for each month and the change in productivity from month to month using sales dollars as the measure of output.

Calculate the productivity for each month and the change in productivity from month to month using number of projects as the measure of output. (Round your answers to 4 decimal places. Negative amounts should be indicated by a minus sign.)

1.
Month Labor Expense Number of Projects Productivity Monthly Change
1 $21,310 44
2 $28,800 60
3 $20,392 60
4 $14,194 15
5 $14,926 20
6 $21,592 60

Calculate the productivity for each month and the change in productivity from month to month using sales dollars as the measure of output. (Round your answers to 4 decimal places. Negative amounts should be indicated by a minus sign.)

2.
Month Labor Expense Sales Dollars Productivity Monthly Change
1 $21,310 $21,090
2 $28,800 $22,900
3 $20,392 $29,780
4 $14,194 $19,560
5 $14,926 $22,680
6 $21,592 $22,520
Answer #1
Answer 1:-
Month Labour Expense(u) Number of Projects(v) Prodcutivity(w) = (v)/(u) Monthly Change Monthly Change(FORMULA)
1           21,310 44 0.0021 0.00000 0
2           28,800 60 0.0021 0.00000 =(G5/G4)-1
3           20,392 60 0.0029 0.41232 =(G6/G5)-1
4           14,194 15 0.0011 -0.64083 =(G7/G6)-1
5           14,926 20 0.0013 0.26794 =(G8/G7)-1
6           21,592 60 0.0028 1.07382 =(G9/G8)-1
Answer 2:-
Month Labour Expense(u) Sales Dollars(v) Prodcutivity(w) = (v)/(u) Monthly Change Monthly Change(FORMULA)
1           21,310              21,090 0.9897 0.0000 0
2           28,800              22,900 0.7951 -0.1966 =(G15/G14)-1
3           20,392              29,780 1.4604 0.8366 =(G16/G15)-1
4           14,194              19,560 1.3780 -0.0564 =(G17/G16)-1
5           14,926              22,680 1.5195 0.1026 =(G18/G17)-1
6           21,592              22,520 1.0430 -0.3136 =(G19/G18)-1